OpenStudy (anonymous):

How do I factor this equation: a^2+10ab-39b^2

OpenStudy (anonymous):

What factors of -39 add up to 10? \(13\times -3=-39\), and \(13+(-3)=10\). So, you can factor the expression as \[a^2+10ab-39b^2=(a+13b)(a-3b)\]